# Data-Adaptive Identification of Subpopulations Vulnerable to Chemical Exposures using Stochastic Interventions

### v1

@@ 3.4 Cross-Validation
As discussed, sample splitting is required to avoid bias incurred by discovering subregions of the covariate space and estimating the effect of subpopulation intervention using the same data. For this we use K-fold cross-validation in which the sample data are split into K-folds of roughly equal size. For example, consider $K$ = 10, and define $k=1$. The $K-1/K$ portion of the sample is used to define the effect modification parameter of interest (here 90%) and the $1/K$ (10%) is used to estimate the parameter, the subpopulation intervention effect. Let $T_{n,k}\subset\{1,\ldots,n\}$ denote the observations indices in the k-th training sample and $V_{n,k}$ denote the observations indices in the k-th validation sample for $k=1,\ldots,K$. The following paragraphs describe how through cross-validation the $T_{n,k}$ (training) portion of the data is used to identify the maximal exposure-effect modifier sets in the mixture and $V_{n,k}$ (validation) is used to estimate our subpopulation stochastic shift intervention parameter.
Section 3.3 described our approach to discovering the covariate subregion with the maximum difference in the intervention effect, this is done using the $T_{n,k}$ training data. More specifically, we train our Q and g estimators using this training data and get targeted estimates for the IIE and ICE. These are then regressed onto our t-statistic tree algorithm to find the region in the covariate space with the largest significant difference in average change due to intervention. The result of this process is a rule such as Sex $=$ 1 or Age $>$ 52 and BMI $>$ 20, etc. which is the region that has the maximum effect difference compared to the complement to that region found in the training estimates.
Now, treating this rule as fixed, we get targeted predictions for our IIE and ICE in the validation sample ($V_{n,k}$) from the Q, g, and targeting models trained in the training fold. We then subset the validation data based on the maximum effect modifying rule found in training and simply take the average of the validation IIE to get point estimates and calculate standard errors from the ICE in the subregions indicated by the rule.
This process is done for each fold of the data and therefore may result in different regions found to be the maximum effect modifying region set.
@@ 3.5 Pooled TMLE
Upon completion of the cross-estimation procedure, a pooled TMLE update provides a summary measure of the oracle maximum effect modification parameter across k-folds. In each fold, we have estimates for Q and g for the respective regions found to be the maximum effect modifying region and it’s complement (again, these may vary). For the pooled TMLE update we stack the estimates for these nuisance parameters from the estimation samples and run a pooled TMLE update on the cumulative initial estimates. The resulting average is then used for parameter estimation. The pooled estimate across the folds can be succinctly represented as:
$\theta=\frac{1}{K}\sum_{k}\Psi_{F_{T_{n,k}}}\left(V_{n,k}\right)(P_{0})$ (17)
Where $\Psi_{F_{T_{n,k}}}$ identifies optimal exposure-covariate modifier set and subsequently formulates a plug-in estimator using training data $T_{n,k}$. $V_{n,k}$ refers to the estimation sample of the data for deriving our estimates. Essentially, this provides an average estimate for the regions $v$ and $v^{c}$ found to have the maximum intervention effect difference across the folds. If $v$ and $v^{c}$ correspond to the same region across all the folds, such as consistently finding age $<22$ for $v$ and age >=>=> = 22 for $v^{c}$ then these pooled estimates are interpretable and provide more power in the regional estimates as they leverage the full data, resulting in tighter confidence intervals.
@@ 3.6 Pooled TMLE for Marginal Exposure Effects
In the case where exposure is a mixture, it is often of interest to the analyst to understand the importance for each exposure relative to the others. Because our approach requires estimation of a shift on each exposure and derivation of the ICE in order to search for the exposure-covariate region with maximum effect modification, we easily can provide CV-TMLE estimates of stochastic shift. That is, for each exposure we simply take our $Q$ and $g$ estimators, trained on the training data, and get the predicitons for the validation data. We then stack our initial estimates $Q$ and $g$, do a pooled TMLE on the full data, and update our counterfactuals of the change in the outcome due to a change in each exposure. This provides estimates for the impact of an exposure change for each exposure with valid inference and can be used to assess for exposure importance.
4 Density Estimation of the Exposures
Estimating the density of each exposure is a crucial component of our methodology to evaluate the effects of the intervention both on average across the population and in subpopulations. We provide two approaches to density estimation, direct density estimation, and a reparameterization based on classification, depending on user input.
4.1 Direct Density Estimation
Super Learner Algorithm: Super Learner operates by training a variety of base learners (individual algorithms) on the data and then combining their predictions using a meta-learner. The meta-learner, in this case, is a non-negative least squares (NNLS) algorithm, which assigns weights to each base learner based on their performance. This approach leverages the strengths of different algorithms to produce a more accurate and robust predictive model.
Homoscedastic Errors (HOSE) Learner: The HOSE learner assumes that the variance of the errors is constant (homoscedasticity) across the range of the predicted values. It employs a semi-parametric approach to density estimation, where the mean of the distribution is estimated using a Super Learner composed of multiple base learners.
Heteroscedastic Errors (HESE) Learner: The HESE learner, in contrast, allows for varying error variances (heteroscedasticity) across different levels of the predicted values. It employs a two-step semi-parametric approach. The mean of the distribution is estimated using a similar Super Learner as in HOSE. An additional Super Learner is used to model the variance of the errors, allowing the error variance to change depending on the value of the predictors.
Although direct estimation is straightforward in principle, it can be challenging in practice due to the scarcity of effective data-adaptive estimators for high-dimensional conditional densities in the machine learning and statistical literature. As such, we also provide another approach.
4.2 Reparameterization of Density Estimation via Classification
As an alternative to direct density estimation, we propose using a novel approach that redefines the problem of estimating the density ratio $H(A,W)$ as a classification problem in an augmented data set [CITE]. This method takes advantage of the fact that classification algorithms are well-developed and widely available in the machine learning literature compared to density estimators.
To apply this approach, we construct an augmented dataset consisting of duplicated observations. For each original observation, we create a duplicate with the observed exposure $A_{t}$ and another with the exposure under the intervention $A^{d}_{t}$. We introduce an indicator variable $\xi$, which is set to 1 if the duplicated observation corresponds to the exposure under intervention, and 0 otherwise. The augmented data set can be represented as:
$\{(W_{\xi,i,t},A_{\xi,i,t},\xi_{\xi,i}):\xi=0,1;i=1,\ldots,n\},$
where $\xi_{\xi,i}=\xi$ indexes the duplicates, $W_{\xi,i,t}=W_{i,t}$ is the history variable, and $A_{\xi,i,t}=\xi\cdot A^{d}_{i,t}+(1-\xi)\cdot A_{i,t}$ is the natural exposure level if $\xi=0$, and the intervened exposure level if $\xi=1$.
We denote the probability distribution of $(W_{t},A_{t},\xi)$ in the augmented dataset by $P_{\xi}$, with corresponding density $p_{\xi}$. We define the parameter $u_{\xi}$ of $P_{\xi}$ as:
$u_{\xi}(a_{t},w_{t})=P_{\xi}(\xi=1|A_{t}=a_{t},W_{t}=w_{t}).$
The density ratio $H(A,W)$ can be expressed in terms of $u_{\xi}$ as follows:
$H_{t}(a_{t},w_{t})=\frac{p_{\xi}(a_{t},w_{t}|\xi=1)}{p_{\xi}(a_{t},w_{t}|\xi=0)}=\frac{P_{\xi}(\xi=1|A_{t}=a_{t},W_{t}=w_{t})}{P_{\xi}(\xi=0|A_{t}=a_{t},W_{t}=w_{t})}=\frac{u_{\xi}(a_{t},w_{t})}{1-u_{\xi}(a_{t},w_{t})},$
where the first equality follows from the definition of $r_{t}$ and the definition of conditional density, the second from Bayes’ rule and the observation that $P_{\xi}(\xi=1)=P_{\xi}(\xi=0)=\frac{1}{2}$, and the third from the definition of $u_{\xi}$.
By estimating $u_{\xi}$ in the augmented dataset using any classification method available in the machine learning and statistical learning literature (e.g., Super Learning), we can indirectly estimate the density ratio. This classification-based approach ensures that the estimator retains desirable properties such as asymptotic normality, provided that appropriate cross-validation and cross-fitting techniques are employed. Furthermore, this reclassification approach is substantially faster computationally which is necessary, especially in our proposed methodology, where the conditional density is needed for each exposure and higher folds in cross-validation provide more consistent results. [CITE]
In our implementation, users can choose between direct density estimation and the reparameterization approach via a user-defined parameter.
5 Summary of Effect Modification Algorithm
Our method of identifying subpopulations differentially impacted by mixed exposures involves several key steps: 1. Identification of Subpopulations: We aim to identify subpopulations that are differentially impacted by mixed exposures. This is achieved by establishing an oracle target parameter that defines the maximal effect modifier relationship. This oracle parameter serves as the benchmark for identifying the subpopulations with the most significant differential impacts from an intervention. 2. Estimation of Oracle Parameter: Once the oracle target parameter is defined, we estimate it by finding the subregions within the covariate space that best approximate our oracle. This involves developing a t-statistic tree, which is a recursive partitioning algorithm that finds the significantly different subregions based on targeted counterfactuals of a shift intervention and influence curve estimates for each exposure. 3. Recursive Partitioning Algorithm: Our recursive partitioning algorithm, guided by t-statistics, identifies the subregions where the impact of the intervention is maximally different. By comparing the averages from the individual intervention effects (IIE) and standard error estimates from the influence curves (IC) in each covariate region, the algorithm splits based on decreasing p-values. 4. Honest Estimation through Cross-Validation: To ensure honest estimation and avoid overfitting, we employ cross-estimation. We use training folds to generate the regions that match our oracle target parameter and estimation samples to estimate the impact of intervention in those regions. This cross-validation process ensures that our results are reliable and generalizable. 5. Efficient Estimation with TMLE: We use TMLE to obtain efficient estimates of the individual intervention effects for each exposure. This method provides robust estimates and valid confidence intervals, leveraging machine learning to flexibly and accurately estimate the necessary components of the model. 6. Pooled Estimates for Robust Inference: After completing the cross-validation procedure, we perform a pooled TMLE update to provide a summary measure of the oracle maximum effect modification parameter across folds. This pooled estimate combines the information from all folds for the regions $v$ and $v^{c}$.
@@ 6.1 Data Generation with Binary Modifiers
The outcome $Y$ was then simulated based on the exposures and covariates, incorporating effectm modification effects:
@@ 6.3 Performance Metrics
To establish the ground truth for the differential impacts of exposures, we applied a shift intervention to reduce $A_{1}$ by a fixed amount ($\delta=-0.5$) and recalculated the outcome $Y$ to obtain $Y_{\delta}$. The difference between the shifted and unshifted outcomes ($\Psi=Y_{\delta}-Y$) was aggregated to estimate the true average effect within each level of the covariate $W_{3}$, where the true effect modification was created. Here the true average in level 0 for $W_{3}$ is region $V$ and level 1 is $V^{c}$.
@@ 7.3 Continuous Effect Modifier Simulation Results
7.4 EffectXshift Applied to NHANES Dataset
The 2001-2002 National Health and Nutrition Examination Survey (NHANES) cycle provides a robust data set for our analysis, widely recognized for its credibility in the public health domain. This data set includes interviews with 11,039 individuals, of whom 4,260 provided blood samples and consented to DNA analysis. For our analysis, we focused on a subset of 1,007 participants, ensuring complete exposure data. This is comparable to the subset used by Mitro et al. Mitro et al. (2016), who investigated the association between persistent exposure to organic pollutants (POPs) binding to the aryl hydrocarbon receptor (AhR) and leukocyte telomere length (LTL). This data set was also used by Gibson et a., Gibson et al. (2019) who applied various mixture methods to this data. For these reasons, this data set works nicely for comparison of the marginal variable importance effects given our stochastic shift interventions and provides evaluation of effect modification which has otherwise not been assessed for in these methods.
@@ 7.6 Identification of Maximum Effect Modifiers
LBXPCBLA 0.7047676 0.04008078 0.6262 0.7833 age_cent $\leq$ 14 1
LBXPCBLA 0.5625930 0.08915217 0.3879 0.7373 age_cent $>$ 14 1
LBXPCBLA 0.8017565 0.03205413 0.7389 0.8646 age_cent $\leq$ 11 2
LBXPCBLA 0.7246825 0.08895869 0.5503 0.8990 age_cent $>$ 11 2
LBXPCBLA 0.3222981 0.02627014 0.2708 0.3738 age_cent $\leq$ 11 3
LBXPCBLA 0.3271666 0.07242882 0.1852 0.4691 age_cent $>$ 11 3
LBXPCBLA -0.6316206 0.02034907 -0.6715 -0.5917 age_cent $\leq$ 11 4
LBXPCBLA -0.3153549 0.03551484 -0.3850 -0.2457 age_cent $>$ 11 4
LBXPCBLA 0.8372697 0.02145195 0.7952 0.8793 age_cent $\leq$ 14 5
LBXPCBLA 0.6842051 0.07540448 0.5364 0.8320 age_cent $>$ 14 5
LBXPCBLA 0.7044444 0.02454143 0.6563 0.7525 age_cent $\leq$ 11 6
LBXPCBLA 0.6348496 0.06230732 0.5127 0.7570 age_cent $>$ 11 6
LBXPCBLA 0.7364454 0.02360700 0.6902 0.7827 age_cent $\leq$ 17 7
LBXPCBLA 0.6057689 0.08581898 0.4376 0.7740 age_cent $>$ 17 7
LBXPCBLA 0.6412914 0.02309030 0.5960 0.6865 age_cent $\leq$ 11 8
LBXPCBLA 0.5364844 0.06194912 0.4151 0.6579 age_cent $>$ 11 8
LBXPCBLA 0.7782908 0.01962983 0.7398 0.8168 age_cent $\leq$ 11 9
LBXPCBLA 0.5685745 0.07544209 0.4207 0.7164 age_cent $>$ 11 9
LBXPCBLA 0.2118344 0.01881629 0.1750 0.2487 age_cent $\leq$ 17 10
LBXPCBLA 0.3519853 0.08829586 0.1789 0.5250 age_cent $>$ 17 10

### v2

@@ 3.4 Cross-Validation
As discussed, sample splitting is required to avoid bias incurred by discovering subregions of the covariate space and estimating the effect of subpopulation intervention using the same data. To achieve this, we employ $K$-fold cross-validation, where the sample data are partitioned into $K$ folds of approximately equal size. For example, consider $K=10$. In each iteration $k=1,\ldots,K$:
  1. 1.
    Training and Validation Split:
    • The training set $T_{n,k}\subset\{1,\ldots,n\}$ consists of the indices from all folds except the $k$-th fold, comprising $\frac{K-1}{K}$ of the data (e.g., 90% for $K=10$).
    • The validation set $V_{n,k}\subset\{1,\ldots,n\}$ consists of the indices in the $k$-th fold, comprising $\frac{1}{K}$ of the data (e.g., 10% for $K=10$).
  2. 2.
    Stage 1: Subpopulation Identification and Parameter Estimation on Training Set:
    • Use the training data $T_{n,k}$ to identify the maximal exposure-effect modifier sets within the mixture using the recursive partitioning algorithm described in Section 3.3.
    • Estimate the nuisance parameters $\hat{\overline{Q}}$ and $\hat{g}$ using appropriate machine learning methods.
    • Compute the individual intervention effects (IIE) and influence curve estimates (ICE) for each observation in $T_{n,k}$.
  3. 3.
    Stage 2: Parameter Estimation on Validation Set:
    • Apply the subpopulation rule $\mathcal{V}_{k}$ identified from the training set to the validation data $V_{n,k}$, assigning observations to $\mathcal{V}_{k}$ or $\mathcal{V}_{k}^{c}$.
    • Estimate the subpopulation stochastic shift intervention parameter $\Psi_{\mathcal{V}_{k}}(P)$ using the CV-TMLE approach with the nuisance parameters estimated from $T_{n,k}$.
After completing all $K$ folds, we aggregate the estimates across folds to obtain the final estimate of the effect modification parameter:
$\hat{\Psi}_{\text{pooled}}=\frac{1}{K}\sum_{k=1}^{K}\hat{\Psi}_{\mathcal{V}_{k}}.$ (14)
If the identified subpopulation $\mathcal{V}_{k}$ is consistent across folds, this pooled estimate leverages the full data for increased efficiency.
@@ 3.5 Pooled TMLE
Upon completion of the cross-estimation procedure, a pooled TMLE update provides a summary measure of the oracle maximum effect modification parameter across $K$-folds. In each fold $k$, we obtain estimates for $\overline{Q}$ and $g$ for the respective regions identified as the maximum effect-modifying regions and their complements. These regions may vary across different folds.
For the pooled TMLE update, we perform the following steps:
  1. 1.
    Stacking Nuisance Parameters: Stack the estimates of the nuisance parameters $\hat{\overline{Q}}$ and $\hat{g}$ from each training sample $T_{n,k}$.
  2. 2.
    Pooled TMLE Update: Apply a pooled TMLE update on the cumulative initial estimates. This involves:
    • Combining the initial outcome regression estimates and propensity score estimates across all folds.
    • Performing a TMLE update on this pooled data to refine the estimates, ensuring they are both efficient and unbiased.
  3. 3.
    Aggregating Estimates: Compute the pooled estimate of the effect modification parameter by averaging the estimates across all folds:
    $\Psi_{\text{pooled}}=\frac{1}{K}\sum_{k=1}^{K}\Psi_{F_{T_{n,k}}}\left(V_{n,k}\right),$ (15)
    where $\Psi_{F_{T_{n,k}}}$ identifies the optimal exposure-covariate modifier set and formulates a plug-in estimator using the training data $T_{n,k}$. $V_{n,k}$ refers to the validation sample for deriving the estimates.
  4. 4.
    Interpretation of Pooled Estimates: The pooled estimate $\Psi_{\text{pooled}}$ represents an average across folds. If the identified subpopulations $\mathcal{V}_{k}$ and $\mathcal{V}_{k}^{c}$ are consistent across all folds (e.g., consistently identifying age $<22$ for $\mathcal{V}$ and age $\geq 22$ for $\mathcal{V}^{c}$), the pooled estimates are interpretable and provide increased statistical power. This leverages the full data, resulting in tighter confidence intervals and more reliable inference.
@@ 3.6 Pooled TMLE for Marginal Exposure Effects
When analyzing a mixture of multiple exposures, it is often important to assess the relative importance of each exposure component. Our approach facilitates this by estimating a stochastic shift for each exposure and deriving the corresponding influence curve estimates (ICE) to identify regions with maximal effect modification.
Since our target parameter requires stochastic shift interventions for each exposure as a preliminary step before recursive partitioning, we can efficiently leverage these estimates to obtain CV-TMLE for each exposure. The procedure is as follows:
Utilizing Existing Estimates: In the first stage of our methodology, we estimate stochastic shift interventions for each exposure $A_{i}$ in the mixture $\boldsymbol{A}$ and derive the corresponding ICEs. These estimates are obtained through the training phase of our cross-validation process and are applied to the validation folds.
Pooled CV-TMLE Estimates: After estimating the stochastic shifts and ICEs for each exposure across all folds, we aggregate these results to compute CV-TMLE estimates for each exposure shift. This pooling process involves:
  • Saving Validation Estimates: We store the TMLE-updated estimates of the outcome changes due to shifts in each exposure from the validation folds.
  • Aggregating Estimates: By averaging these estimates across all folds, we obtain pooled CV-TMLE estimates for each exposure $A_{i}$. This aggregation enhances the reliability and statistical power of our estimates.
Variable Importance Measures: The pooled CV-TMLE estimates for each exposure shift allow us to assess the relative importance of each exposure within the mixture. Specifically, these estimates quantify the impact of shifting each exposure $\delta_{i}$ on the outcome, providing a clear measure of variable importance. This comprehensive estimation framework not only identifies regions with maximal effect modification but also offers insights into the marginal effects of each exposure, thereby supporting informed and targeted public health interventions.
In summary, because our target parameter inherently requires stochastic shift interventions for each exposure as part of the initial estimation process, we can directly utilize these results to obtain CV-TMLE estimates. This approach simplifies the workflow and enables the derivation of variable importance measures for the entire exposure mixture.
3.7 Exposure Density Estimation
Accurate estimation of the conditional exposure densities $g(A\mid W)$ is crucial for evaluating the effects of stochastic interventions within a causal inference framework. We employ two approaches for density estimation: direct conditional density estimation and a reparameterization strategy that transforms density ratio estimation into a classification problem. Both methods are implemented using the Super Learner ensemble van der Laan et al. (2007), which optimally combines multiple algorithms to produce robust and flexible estimates.
3.7.1 Direct Conditional Density Estimation
The direct approach involves estimating the conditional density $g(A\mid W)$ directly using a Super Learner composed of a diverse library of density estimation algorithms. This method leverages the strengths of various learners to capture complex relationships between exposures and covariates. However, direct density estimation can be challenging in high-dimensional or intricate settings due to the computational complexity and lack of models to estimate $g(A\mid W)$ accurately.
3.7.2 Reparameterization via Classification
To address the limitations of direct density estimation, we implement a reparameterization strategy inspired by the work of Díaz et al. (2023) and on modified treatment policies (MTPs). This approach reframes the problem of estimating the density ratio $\frac{g(A-\delta\mid W)}{g(A\mid W)}$ as a binary classification task. The procedure involves the following steps:
  1. 1.
    Augmented Dataset Construction: Create an augmented dataset by duplicating the original data. In the duplicated subset, apply the stochastic shift intervention $\delta$ to the exposures $A$, while keeping the exposures in the original subset unchanged. Introduce a binary indicator variable $Z$, where $Z=1$ denotes shifted exposures and $Z=0$ denotes unshifted exposures.
  2. 2.
    Classification Model Training: Fit a Super Learner classifier to predict the probability $P(Z=1\mid A,W)$ using the augmented dataset. This model estimates the likelihood that an observation has been subjected to the shift intervention.
  3. 3.
    Density Ratio Estimation: Utilize Bayes’ theorem to transform the predicted probabilities into the desired density ratios:
    $\frac{g(A-\delta\mid W)}{g(A\mid W)}=\frac{P(Z=1\mid A-\delta,W)}{1-P(Z=1\mid A,W)}.$
This reparameterization offers several key advantages:
  • Computational Efficiency and Scalability: By converting density ratio estimation into a classification problem, we can leverage highly optimized classification algorithms within the Super Learner framework, enhancing computational efficiency and scalability, especially in high-dimensional settings.
  • Numerical Stability: Estimating density ratios through classification probabilities reduces the risk of extreme ratio estimates, thereby enhancing the numerical stability of our estimators.
  • Flexibility with Multiple Exposures: This approach naturally extends to scenarios involving multiple exposures, facilitating future work to model the effect modification of joint exposure shifts.
Given these benefits, our simulations and empirical applications utilize the reparameterization strategy for density ratio estimation. This choice is motivated by its superior performance in terms of computational efficiency, estimator stability, which is particularly important in our case where the density ratio needs to be calculated for each exposure.
@@ 4.1 Data Generation with Binary Modifiers
The outcome $Y$ was then simulated based on the exposures and covariates, incorporating effect modification effects:
@@ 4.3 Performance Metrics
To establish the ground truth for the differential impacts of exposures, we applied a shift intervention to reduce $A_{1}$ by 0.5 units, i.e., we set $A_{1}\leftarrow A_{1}-0.5$, and recalculated the outcome $Y$ to obtain $Y_{\delta}$. Other exposures remained unchanged. The difference between the shifted and unshifted outcomes ($\Psi=Y_{\delta}-Y$) was aggregated to estimate the true average effect within each level of the covariate $W_{3}$, where the true effect modification was created. Here the true average in level 0 for $W_{3}$ is region $\mathcal{V}$ and level 1 is $\mathcal{V}^{c}$.
@@ 5.3 Continuous Effect Modifier Simulation Results
6 EffectXshift Applied to NHANES Dataset
The 2001-2002 National Health and Nutrition Examination Survey (NHANES) cycle provides a robust data set for our analysis, widely recognized for its credibility in the public health domain. This data set includes interviews with 11,039 individuals, of whom 4,260 provided blood samples and consented to DNA analysis. For our analysis, we focused on a subset of 1,007 participants, ensuring complete exposure data. This is comparable to the subset used by Mitro et al. Mitro et al. (2016), who investigated the association between persistent exposure to organic pollutants (POPs) binding to the aryl hydrocarbon receptor (AhR) and leukocyte telomere length (LTL). This data set was also used by Gibson et al., Gibson et al. (2019) who applied various mixture methods to this data. For these reasons, this data set works nicely for comparison of the marginal variable importance effects given our stochastic shift interventions and provides evaluation of effect modification which has otherwise not been assessed for in these methods.
@@ 6.2 Identification of Maximum Effect Modifiers
LBXPCBLA 0.7047676 0.04008078 0.6262 0.7833 age $\leq$ 14 1
LBXPCBLA 0.5625930 0.08915217 0.3879 0.7373 age $>$ 14 1
LBXPCBLA 0.8017565 0.03205413 0.7389 0.8646 age $\leq$ 11 2
LBXPCBLA 0.7246825 0.08895869 0.5503 0.8990 age $>$ 11 2
LBXPCBLA 0.3222981 0.02627014 0.2708 0.3738 age $\leq$ 11 3
LBXPCBLA 0.3271666 0.07242882 0.1852 0.4691 age $>$ 11 3
LBXPCBLA -0.6316206 0.02034907 -0.6715 -0.5917 age $\leq$ 11 4
LBXPCBLA -0.3153549 0.03551484 -0.3850 -0.2457 age $>$ 11 4
LBXPCBLA 0.8372697 0.02145195 0.7952 0.8793 age $\leq$ 14 5
LBXPCBLA 0.6842051 0.07540448 0.5364 0.8320 age $>$ 14 5
LBXPCBLA 0.7044444 0.02454143 0.6563 0.7525 age $\leq$ 11 6
LBXPCBLA 0.6348496 0.06230732 0.5127 0.7570 age $>$ 11 6
LBXPCBLA 0.7364454 0.02360700 0.6902 0.7827 age $\leq$ 17 7
LBXPCBLA 0.6057689 0.08581898 0.4376 0.7740 age $>$ 17 7
LBXPCBLA 0.6412914 0.02309030 0.5960 0.6865 age $\leq$ 11 8
LBXPCBLA 0.5364844 0.06194912 0.4151 0.6579 age $>$ 11 8
LBXPCBLA 0.7782908 0.01962983 0.7398 0.8168 age $\leq$ 11 9
LBXPCBLA 0.5685745 0.07544209 0.4207 0.7164 age $>$ 11 9
LBXPCBLA 0.2118344 0.01881629 0.1750 0.2487 age $\leq$ 17 10
LBXPCBLA 0.3519853 0.08829586 0.1789 0.5250 age $>$ 17 10